# On Utility Maximization under Multivariate Fake Stationary Affine Volterra Models

### 06 Apr 2026 v2

@@ 3.3 Optimal strategy for the logarithmic utility maximization problem
[…]
We want to solve the Merton logarithmic utility optimization case, i.e. our aim is to find the value function defined in (3.21) for the logarithmic utility function such that
$\mathcal{V}(x_{0},v_{\infty}):=\operatorname*{\,sup}_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\Big[\log(X_{T}^{\alpha})\Big],\qquad X_{0}^{\alpha}:=x_{0}>0,\quad v_{\infty}:=\mathbb{E}[V_{0}];$ (3.83)
@@ 4 Numerical experiments: The fake stationary rough Heston volatility
Figures 3 below also confirm the Remark on Proposition 3.6, that is the claim that $\psi\leq 0$ for every $\rho_{i}\in(0,1)$ and every $i\in\{1,2\}$ in the exponential utility case.
@@
Plots show that, in addition to the impact of the roughness of asset volatility on the optimal allocation documented in Han and Wong (2020); Aichinger and Desmettre (2021), the stabilizing function also affects the optimal portfolio allocation, and hence the hedging demand over time. The form of the control is consistent across different time scales. The roughness dominates the very short end, while the stabilizing function dictates the long-term behaviour. While the rough volatility models capture high-frequency dynamics with remarkable accuracy, their lack of stationarity make them unsuitable for long-term investment analysis. This framework is then also essential for formulating a well-posed infinite-horizon investment problem, time-series analysis and long-term econometric estimation.
Our illustrations also show that the optimal allocation in the exponential utility case is moderated as the paths of the volatility become rougher. Since $\varsigma^{i}(0)\psi^{i}(T)=\varsigma^{i}(T)\psi^{i}(0)=0$ for $i\in\{1,2\}$, it follows that, in the power utility case, the vector of amounts invested in each stock coincides at the beginning and at the end of the time horizon, as illustrated in the left panels of Figures 4.
@@ 5 Proofs of the main results
Let $W$, $W^{\top}$ be two independent $d$-dimensional Wiener processes and $V$ a weak solution to the inhomogeneous Volterra Square root Equation (2.2). Let $g_{1}$ and $g_{2}$ two deterministic processes with values in $\mathbb{R}^{d}$ such that $g_{1,i},g_{2,i}\in L^{\infty}(\mathbb{R}^{+},\mathbb{R})$ for $1\leq i\leq d$. Then the local martingale
$Z_{t}:=\mathcal{E}(\int_{0}^{t}\sum\limits_{i=1}^{d}g_{1,i}(s)\sqrt{V_{s}^{i}}dW^{i}_{s}+\int_{0}^{t}\sum\limits_{i=1}^{d}g_{2,i}(s)\sqrt{V_{s}^{i}}dW^{\top,i}_{s})$
@@
Lemma 5.2 (Martingale property).
Let $g$ be a deterministic process with values in $\mathbb{R}$ and bounded such that $g\in L^{\infty}([0,T],\mathbb{R})$. Let us denote
${Z}_{t}:=\mathcal{E}(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s})=\exp\Big(\int_{0}^{t}g(s)\alpha_{s}^{\top}\,dB_{s}-\frac{1}{2}\int_{0}^{t}g^{2}(s)\left|\alpha_{s}\right|^{2}ds\Big),\quad t\in[0,T].\vskip-5.69046pt$ (5.86)
where $(\alpha_{t})_{t\in[0,T]}$ is a $\mathbb{R}^{d}$- valued process belonging to $L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})$}. Then $(Z_{t})_{t\in[0,T]}$ is a martingale.
Proof: This is a direct consequence of the fact that $g\in L^{\infty}([0,T],\mathbb{R})$ and $\alpha\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ together with Novikov’s condition.∎
@@ 2. $(b)$
where $F_{T}^{\alpha}$ is an $(\mathbb{F},\mathbb{P})$-local martingale as $(\alpha,\Lambda)$ are in $L^{2,{loc}}_{\mathbb{F}}([0,T])$ (Dooléans-Dade exponential) and write
$F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}\gamma\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\Big(\gamma^{2}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}\Big)ds+\gamma\int_{0}^{t}\alpha_{s}^{\top}dB_{s}+\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
At this stage, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $g_{1,i}(t)=\frac{\gamma}{1-\gamma}\rho_{i}\theta_{i}+(1+\frac{\gamma}{1-\gamma}\rho_{i}^{2})\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\in L^{\infty}([0,T],\mathbb{R})$ and $g_{2,i}(t)=\frac{\gamma}{1-\gamma}\sqrt{1-\rho_{i}^{2}}\big(\theta_{i}+\rho_{i}\nu_{i}\varsigma^{i}(t)\psi^{i}(T-t)\big)\in L^{\infty}([0,T],\mathbb{R})$ for $1\leq i\leq d$. Consequently, inserting the candidate (3.50) for the optimal strategy $\alpha^{*}$ in the above Equation (5.94), and observing that $D_{s}(\alpha_{s}^{*})=0\;\forall\,s\in[0,T]$, then taking the expectation, it is straightforward that:
@@
where the last inequality comes from the martingality of $F_{T}^{\alpha^{*}}$ and the desired result about the first part of the proof (equality (1)) is completed. It remains to show the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$.
@@
Since $\alpha_{t}$ is bounded by assumption, we can introduce a new probability measure $\mathbb{Q}$ with Radon-Nikodym density or derivative at $F_{T}$ as
@@
by Lemma 5.2. By Girsanov’s theorem, we may write:
@@
$\qquad+\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{\nu_{i}^{2}}{2}\big(\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}V_{s}^{i}\,ds+\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}\,d\widehat{W}^{i}_{s}\Big).$
@@
$\hskip 79.6678pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}$
$\hskip 170.71652pt+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds\Big)\Big].$ (5.99)
@@
$\hskip 59.75095pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}^{*}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{*i}_{s}\Big)\Big].$
Since $\psi$ is continuous, it is bounded and therefore the stochastic exponential is a true $\mathbb{Q}^{*}$- martingale with expectation $1$ by Lemma 5.1 with $g_{2}=0$ and $g_{1,i}(s)=\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\in L^{\infty}(\mathbb{R}^{+},\mathbb{R})$ . Thus we get
@@
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\sum_{i=1}^{d}\int_{0}^{T}\frac{\nu_{i}^{2}}{2}(\varsigma^{i}(s)\psi^{i}(T-s))^{2}V_{s}^{i}ds+\sum_{i=1}^{d}\int_{0}^{T}\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}d\widehat{W}^{i}_{s}\Big)\Big].$
@@
Proof of Theorem 3.12:
@@
Step 1 (Proof of Admissibility:) We begin by proving the admissibility of the candidate in equation (3.78) for the optimal portfolio strategy $\alpha^{*}$, and, on the way, we deduce the bound in equation (3.81).

### 07 May 2026 v3

@@ 3.3 Optimal strategy for the logarithmic utility maximization problem
$\mathcal{A}=\left\{\alpha=(\alpha_{t})_{t\in[0,T]}\in{L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}\right\}.$ (3.87)
Note that, for any $\alpha\in\mathcal{A}$, Equation (3.32) has a positive solution $(X_{t}^{\alpha})_{t\in[0,T]}$. We want to solve the Merton logarithmic utility optimization case, i.e. our aim is to find the value function defined in (3.23) for the logarithmic utility function such that
$\hskip 28.45274pt\mathcal{V}(x_{0},v_{\infty})$ $=\operatorname*{\,sup}_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\Big[\log(X_{T}^{\alpha})\Big],\qquad X_{0}^{\alpha}=x_{0}>0,\quad v_{\infty}=\mathbb{E}[V_{0}]$ (3.88)
$=\log(x_{0})+\sup_{\alpha(\cdot)\in\mathcal{A}}\mathbb{E}\Bigl[\int_{0}^{T}\bigl(r(s)+\alpha_{s}^{\top}\lambda_{s}-\tfrac{1}{2}\left|\alpha_{s}\right|^{2}\bigr)ds+\int_{0}^{T}\alpha_{s}^{\top}\,dB_{s}\Bigr].$
@@ 4 Numerical experiments:The fake stationary rough Heston volatility
Figures 3 below also confirm the Remark on Proposition 3.6, that is the claim that $\psi\leq 0$ for every $\rho_{i}\in(0,1)$ and every $i\in\{1,2\}$ in the exponential utility case. The lower the Hurst coefficient $H$, the more negative is $\psi$. In the power utility case however, the left panel of the figure shows that, in accordance with Han and Wong (2020), $\psi(t)$ is positive for $t>0$ and its value becomes larger for a smaller hurst exponent $H$.
@@
Plots of Figure 4 below show that, in addition to the impact of the roughness of asset volatility on the optimal allocation documented in Han and Wong (2020); Aichinger and Desmettre (2021), the stabilizing function also affects the optimal portfolio allocation, and hence the hedging demand over time. The form of the control is consistent across different time scales. The roughness dominates the very short end, while the stabilizing function dictates the long-term behaviour.
While the rough volatility models capture high-frequency dynamics with remarkable accuracy, their lack of stationarity make them unsuitable for long-term investment analysis. This framework is then also essential for formulating a well-posed infinite-horizon investment problem, time-series analysis and long-term econometric estimation.
@@
Our illustrations further reveals that, the lower the investor’s risk aversion, the smaller his optimal allocation in the power utility case. Additionally, the vector of amounts invested in each stock coincides at both the beginning and the end of the time horizon, as shown in the left panels of Figures 4 (Since $\varsigma^{i}(0)\psi^{i}(T)=\varsigma^{i}(T)\psi^{i}(0)=0$ for $i\in\{1,2\}$).
In contrast, the exponential utility suggests a completely different perspective: the optimal allocation becomes more moderated as the volatility paths become rougher and risk aversion increases.
Economically, in line with Han and Wong (2020), these results indicate that distinct forms of risk aversion, captured by concave utility functions, display varying degrees of sensitivity in investment demand to the roughness of volatility.
@@ 5 Proofs of the main results
Let $g_{1}$ and $g_{2}$ be two deterministic $\mathbb{R}$-valued bounded processes such that $g_{1},g_{2}\in L^{\infty}([0,T],\mathbb{R})$. Let $\left(\lambda,\Lambda\right)\in{L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})\times L^{2}_{\mathbb{F}}([0,T],\mathbb{R}^{d})}$ be two $\mathbb{R}^{d}$-valued stochastic processes, defined in (2.14) and (3.26) respectively. Let $B$, $W$ be two $d$-dimensional Wiener processes and $\Sigma$ their corellation matrix, then under assumption 3.1, for any $\kappa\geq 0$ the local martingale
$Z_{t}:=\mathcal{E}\Big(\int_{0}^{t}g_{1}(s)(\lambda_{s}+\kappa\Sigma\Lambda_{s})^{\top}dB_{s}+g_{2}(s)\Lambda_{s}^{\top}dW_{s}\Big)$
@@
Proof: Define $M_{t}:=\int_{0}^{t}g_{1}(s)(\lambda_{s}+\kappa\Sigma\Lambda_{s})^{\top}dB_{s}+\int_{0}^{t}g_{2}(s)\Lambda_{s}^{\top}dW_{s}.$ Then $Z=\mathcal{E}(M)$ is a positive local martingale. We compute its quadratic variation. Using $d\langle B,W\rangle_{t}=\Sigma dt$, we obtain
$\langle M\rangle_{t}$ $=\int_{0}^{t}g_{1}(s)^{2}|\lambda_{s}+\kappa\Sigma\Lambda_{s}|^{2}ds+\int_{0}^{t}g_{2}(s)^{2}|\Lambda_{s}|^{2}ds+2\int_{0}^{t}g_{1}(s)g_{2}(s)(\lambda_{s}+\kappa\Sigma\Lambda_{s})^{\top}\Sigma\Lambda_{s}ds$
$\leq\int_{0}^{t}2g_{1}(s)^{2}|\lambda_{s}+\kappa\Sigma\Lambda_{s}|^{2}ds+\int_{0}^{t}g_{2}(s)^{2}\big(1+|\Sigma|^{2}\big)|\Lambda_{s}|^{2}ds\leq 4C\big(1+|\Sigma|^{2}\big)\int_{0}^{t}\left(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\right)ds$
where we use the elementary arithmetic mean-geometric mean (AM-GM) inequality $|ab|\leq(|a|^{2}+|b|^{2})/2$ and set $C:=(1+\kappa^{2})\cdot\max(1,\|g_{1}\|_{\infty,T},\|g_{2}\|_{\infty,T})\geq 1$ owing to the boundedness of $g_{1},g_{2}$. Hence, we have
$\mathbb{E}\!\Big[\exp\!\Big(\tfrac{1}{2}\langle M\rangle_{T}\Big)\Big]\leq\mathbb{E}\Big[\exp\Big({a(2C)\int_{0}^{T}\Big(|\lambda_{s}|^{2}+|\Lambda_{s}|^{2}\Big)ds}\Big)\Big]<\infty,$ (5.91)
thanks to condition (3.29) (since Assumption 3.1 is in force) with constant $a(2C)=2C\big(1+|\Sigma|^{2}\big)$. Consequently, Novikov’s condition holds, so $\mathcal{E}(M)$ is a true martingale.∎
@@ 2. $(b)$
where $F_{t}^{\alpha}$ is an $(\mathbb{F},\mathbb{P})$-local martingale as $(\alpha,\Lambda)$ are in $L^{2,{loc}}_{\mathbb{F}}([0,T])^{2}$ (Dooléans-Dade exponential) and write
$F_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}\gamma\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)=\operatorname{exp}\Big(-\frac{1}{2}\int^{t}_{0}\Big(\gamma^{2}\left|\alpha_{s}\right|^{2}+\left|\Lambda_{s}\right|^{2}+2\gamma\alpha_{s}^{\top}(\Sigma\Lambda_{s})\Big)ds+\gamma\int_{0}^{t}\alpha_{s}^{\top}dB_{s}+\int_{0}^{t}\Lambda_{s}^{\top}dW_{s}\Big).$
At this stage, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $\kappa=1$, $g_{1}(t)=\frac{\gamma}{1-\gamma}$ and $g_{2}(t)=1$ for every $t\in[0,T]$. Consequently, inserting the candidate (3.52) for the optimal strategy $\alpha^{*}$ in the above Equation (5.99), and observing that $D_{s}(\alpha_{s}^{*})=0\;\forall\,s\in[0,T]$, then taking the expectation, it is straightforward that:
@@
where the last inequality comes from the martingality of $F_{T}^{\alpha^{*}}$ and the desired result about the first part of the proof (equality (1)) is completed. It remains to show the inequality (2) for arbitrary admissible portfolio strategies $\alpha\in\mathcal{A}$ defined in (3.53).
@@
Since by assumption, $\alpha\in\mathcal{A}$ redefined in (3.53), we can introduce a new probability measure $\mathbb{Q}$ with Radon-Nikodym density or derivative at $\mathcal{F}_{T}$
@@
thanks to Novikov’s condition. (In fact, since $\gamma\in(0,1)$, the function $x\mapsto x^{\gamma^{2}}$ is concave. Jensen’s inequality together with (3.53) implies that $\mathbb{E}\big[\exp\big(\frac{\gamma^{2}}{2}\int_{0}^{T}\left|\alpha_{s}\right|^{2}ds\big)\big]<\infty$ so that $(Z_{t})_{t\in[0,T]}$ is a martingale.) By Girsanov’s theorem, we may write:
@@
$\hskip 0.56917pt\qquad+\sum_{i=1}^{d}\int_{0}^{T}\Big(-\frac{\nu_{i}^{2}}{2}\big(\varsigma^{i}(s)\psi^{i}(T-s)\big)^{2}V_{s}^{i}\,ds+\nu_{i}\varsigma^{i}(s)\psi^{i}(T-s)\sqrt{V_{s}^{i}}\,d\widehat{W}^{i}_{s}\Big)=\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds$
$+\sum_{i=1}^{d}\int_{0}^{T}\Big(\frac{\gamma}{2(1-\gamma)}\theta_{i}^{2}+F_{i}(s,\psi(T-s))\Big)g_{0}^{i}(s)\,ds-\frac{1}{2}\int_{0}^{T}|\Lambda_{s}|^{2}\,ds+\int_{0}^{T}\Lambda_{s}^{\top}d\widehat{W}^{i}_{s}.$
@@
$\hskip 56.9055pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\frac{1}{2}\int_{0}^{T}|\Lambda_{s}|^{2}\,ds+\int_{0}^{T}\Lambda_{s}^{\top}d\widehat{W}^{i}_{s}+\sum_{i=1}^{d}\int_{0}^{T}\frac{\gamma(\gamma-1)}{2}\hat{\pi}_{s,i}^{2}V^{i}_{s}\,ds\Big)\Big].$ (5.104)
@@
$\hskip 102.43008pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}^{*}}\Big[\exp\Big(-\frac{1}{2}\int_{0}^{T}|\Lambda_{s}|^{2}\,ds+\int_{0}^{T}\Lambda_{s}^{\top}d\widehat{W}^{*i}_{s}\Big)\Big].$
Note that, the stochastic exponential is a true $\mathbb{Q}^{*}$- martingale with expectation $1$ by Lemma 5.1 with $g_{1}=0$ and $g_{2}(s)=1$ for every $s\in[0,T]$. Consequently, we get
@@
$\hskip 62.59596pt\times\mathbb{E}_{x_{0},V_{0}}^{\mathbb{Q}}\Big[\exp\Big(-\frac{1}{2}\int_{0}^{T}|\Lambda_{s}|^{2}\,ds+\int_{0}^{T}\Lambda_{s}^{\top}d\widehat{W}^{i}_{s}\Big)\Big].$
@@
Proof of Theorem 3.12 Step 1 (Proof of the Martingale optimality principle for $J^{\alpha}$.) We show that $J_{t}^{\alpha}$ in (3.86) fulfills the martingale optimality principle 3.2. For the first condition, note that $\Gamma_{T}=1$ and hence $J_{T}^{\alpha}=-\frac{1}{\gamma}\exp\left(-\gamma X^{\alpha}_{T}\right)=U(X^{\alpha}_{T})$. Since $\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$, $J^{\alpha}_{0}=-\frac{1}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)\Gamma_{0}$ is a constant independent of $\alpha\in{\mathcal{A}}$ and thus the second condition is also satisfied. In order to show that the third condition is also fulfilled, we apply Itô’s formula on $J_{t}^{\alpha}$ defined in (3.86). Let’s us consider the Problem (3.60) with an arbitrary strategy $\alpha\in\mathcal{A}$. To ease notations, we set $U_{t}:=-\gamma e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}+\log(\Gamma_{t})$ and $h_{t}:=\lambda_{t}+\Sigma\Lambda_{t}$. Note that
$J^{\alpha}_{t}:=-\frac{1}{\gamma}\exp(U_{t})\quad\text{and}\quad dJ_{t}^{\alpha}=J_{t}^{\alpha}\Big(dU_{t}+\frac{1}{2}d\langle U\rangle_{t}\Big)$
For any admissible strategy $\alpha\in\mathcal{A}$, we write by Itô’s lemma:
$dU_{t}$ $=\Big(\gamma r(s)e^{\int_{t}^{T}r(s)ds}X^{\alpha}_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\big(r(t)X^{\alpha}_{t}+\alpha_{t}^{\top}\lambda_{t}\big)\Big)dt-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+d\log(\Gamma_{t})$
$=\big(-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}\lambda_{t}+f(Y_{t},\Lambda_{t},U_{t})-\frac{1}{2}\left|\Lambda_{t}\right|^{2}\big)dt+\Lambda_{t}^{\top}dW_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}$
where $f$ is the driver or drift term in Equation (3.75). Consequently, by Itô’s product rule, combined with the property of $\Gamma$ in Equation (3.75) of Proposition 3.11, one may write
$dJ_{t}^{\alpha}$ $=\;J_{t}^{\alpha}\big(-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}\lambda_{t}+f(Y_{t},\Lambda_{t},U_{t})-\frac{1}{2}\left|\Lambda_{t}\right|^{2}+\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha^{\top}_{t}\alpha_{t}-\gamma e^{\int_{t}^{T}r(s)ds}\alpha^{\top}_{t}\big(\Sigma\Lambda_{t}\big)$
$+\frac{1}{2}\left|\Lambda_{t}\right|^{2}\big)dt+J_{t}^{\alpha}\big(-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+\Lambda_{t}^{\top}dW_{t}\big)=J_{t}^{\alpha}\Big(D_{t}(\alpha_{t})dt-\gamma e^{\int_{t}^{T}r(s)ds}\alpha_{t}^{\top}dB_{t}+\Lambda_{t}^{\top}dW_{t}\Big).$
where the drift factor takes the form:
$D_{t}(\alpha)=$ $\frac{\gamma^{2}}{2}e^{\int_{t}^{T}2r(s)ds}\alpha^{\top}\alpha-\gamma e^{\int_{t}^{T}r(s)ds}\alpha^{\top}h_{t}+\frac{1}{2}h_{t}^{\top}h_{t}.$
Differentiating $D_{t}(\alpha)$ with respect to $\alpha$ and checking the second order condition, the maximizer $\alpha^{*}_{t}:=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}h_{t}=\frac{1}{\gamma}e^{-\int_{t}^{T}r(s)ds}\left(\lambda_{t}+\Sigma\Lambda_{t}\right)$ for every $t\in[0,T]$ that is the strategy given by Equation (3.82). Evaluating the drift factor $D_{t}$ at $\alpha^{*}_{t}$ show that $D_{t}(\alpha^{*}_{t})$ vanishes to 0. Note $D_{t}(\alpha)$ is a quadratic function on $\alpha$ and $\gamma-1<0$. As $D_{t}(\alpha^{*}_{t})=0$, then $D_{t}(\alpha_{t})\geq 0$ for any admissible strategy $\alpha$. Moreover, solving the stochastic differential equation for $J_{t}^{\alpha}$ yields
$\hskip 56.9055ptJ_{t}^{\alpha}=-\frac{\Gamma_{0}}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}F_{t}^{\alpha}$ (5.105)
$\text{where}\hskip 113.81102ptF_{t}^{\alpha}=\mathcal{E}\Big(\int_{0}^{t}-\gamma e^{\int_{s}^{T}r(u)du}\alpha_{s}^{\top}dB_{s}+\Lambda_{s}^{\top}dW_{s}\Big)$
Following our assumptions on the admissible strategies in Definition 3.8 and Proposition 3.11, we have that $\left(\alpha,\Lambda\right)\in{L^{2,loc}_{\mathbb{F}}([0,T],\mathbb{R}^{d})^{2}}$ and thus, the stochastic exponential $F_{t}^{\alpha}$ is a $(\mathbb{F},\mathbb{P})$-local martingale (which follows from the basic properties of the Dooléans-Dade exponential). Therefore, there exists a sequence of stopping times $\{\tau_{n}\}_{n\geq 1}$ satisfying $\lim_{n\rightarrow\infty}\tau_{n}=T$, $\mathbb{P}$-a.s., such that that $F^{\alpha}_{t\wedge\tau_{n}}$ is a positive martingale for every $n$.
Furthermore, $\frac{\Gamma_{0}}{\gamma}\exp\big(-\gamma e^{\int^{T}_{0}r(u)du}x_{0}\big)e^{\int^{t}_{0}D_{s}(\alpha_{s})ds}$ is non-increasing. Therefore, $J^{\alpha}_{t\wedge\tau_{n}}$ is a supermartingale. Then for $s\leq t$, $\mathbb{E}[J^{\alpha}_{t\wedge\tau_{n}}|{\mathcal{F}}_{s}]\leq J^{\alpha}_{s\wedge\tau_{n}}$. It implies that for any set $A\in{\mathcal{F}}_{s}$,
for every $n$. Since $\left(\exp\big(-\gamma e^{\int^{T}_{t\wedge\tau_{n}}r(u)du}X_{t\wedge\tau_{n}}\big)\right)_{n\geq 0}$ is uniformly integrable (Definition 3.8 $(c)$) and $\Gamma$ is bounded (see the last claim of Proposition 3.11), $\left(J^{\alpha}_{t\wedge\tau_{n}}\right)_{n}$ and $\left(J^{\alpha}_{s\wedge\tau_{n}}\right)_{n}$ are uniformly integrable. Let $n\rightarrow\infty$, then $\mathbb{E}[J^{\alpha}_{t}\mathds{1}_{A}]\leq\mathbb{E}[J^{\alpha}_{s}\mathds{1}_{A}]$. Consequently, we deduce that $J^{\alpha}$ is a supermartingale for every arbitrary admissible strategy $\alpha$. It remains to show that $J_{t}^{\alpha^{*}}$ is a true martingale for the optimal strategy $\alpha^{*}$ in which case $e^{\int_{0}^{t}D_{s}(\alpha_{s}^{*})ds}=1$ and therefore the above equation (5.105), when inserting the candidate (3.82) for the optimal strategy $\alpha^{*}$ reads $J_{t}^{\alpha^{*}}=-\frac{\exp(Y_{0})}{\gamma}\exp\left(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\right)F_{t}^{\alpha^{*}}$. Now, for $\alpha_{t}=\alpha^{*}_{t}$, $F_{t}^{\alpha^{*}}$ is a $\mathbb{P}$-martingale with expectation $1$ by Lemma 5.1 with $g_{1}(t)=-1$ and $g_{2}(t)=1$ for every $t\in[0,T]$. Subsequently, $J^{\alpha^{*}}_{t}$ is a true martingale and taking the expectation, its boils down that:
$\mathbb{E}_{x_{0},V_{0}}{\big[-\frac{1}{\gamma}\exp\big(-\gamma X^{\alpha^{*}}_{T}\big)\big]}$ $:=\mathbb{E}_{x_{0},V_{0}}{\big[J_{T}^{\alpha^{*}}\big]}=-\frac{\Gamma_{0}}{\gamma}\exp\Big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\Big)\mathbb{E}_{x_{0},V_{0}}{\big[F_{T}^{\alpha^{*}}\big]}$
$=-\frac{1}{\gamma}\exp\Big(-\gamma e^{\int_{0}^{T}r(s)ds}x_{0}\Big)\Gamma_{0}.$
where the last inequality comes from the fact that $F_{T}^{\alpha^{*}}$ is a true martingale. This completes the proof of the equality (3.85). We have verified all conditions required by martingale optimality principle in Definition 3.2, except for the admissibility of $\alpha^{*}$, which follows from Step 2 below.
@@
Step 2 (Proof of Admissibility:) We now prove the admissibility of the candidate in equation (3.81) for the optimal portfolio strategy $\alpha^{*}$, and, on the way, we deduce the bound in equation (3.84).